At year-end 2016, Wallace Landscaping’s total assets were $1.6 million, and its accounts payable were $345,000. Sales, which in 2016 were $2.7 million, are expected to increase by 30% in 2017. Total assets and accounts payable are proportional to sales, and that relationship will be maintained. Wallace typically uses no current liabilities other than accounts payable. Common stock amounted to $365,000 in 2016, and retained earnings were $300,000. Wallace has arranged to sell $150,000 of new common stock in 2017 to meet some of its financing needs. The remainder of its financing needs will be met by issuing new long-term debt at the end of 2017. (Because the debt is added at the end of the year, there will be no additional interest expense due to the new debt.) Its net profit margin on sales is 6%, and 35% of earnings will be paid out as dividends.

What was Wallace's total long-term debt in 2016? Do not round intermediate calculations. Round your answer to the nearest dollar.

What were Wallace's total liabilities in 2016? Do not round intermediate calculations. Round your answer to the nearest dollar. How much new long-term debt financing will be needed in 2017? (Hint: AFN - New stock = New long-term debt.) Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

A)Wallace's total liabilities in 2016

Asset= Short term liabilities+long term liabilities+equity

Long term liabilities= Asset-short term liabilities-equity

= $1,600,000 - $345,000 - ($365,000 + $300,000)

= $590,000

Total liabilities = $590,000 + $345,000

= $935,000

B)

Since sales is going to increase in 2017 by 30% and asset and account payable are propotionate to sales.

Hence asset and account payable will also increase by 30%.

So asset in 2017 = $1.6*130%= $2.08 Million

accounts payable in 2017 = $345,000*130%= $448,500

Total Sales in 2016 = $2.7 million * 130 %= $3.51 Million

Since profit margin = 6%

Net income= = $3.51 million * 6% = $210,600

And 35% is dividend so income after dividend = $210,600 * (1- 35%) = $136,890

So in retained earning = beginning retained earning +income after div paid

= $300,000+$136,890 = $436,890

So now we again use the same formula in 2016

long term liabilities= Asset-short term liabilities-equity

= $2,080,000 - $448,500 - $436,890 - $365,000 - $150,000

=$679,610

He already had long term liabilities= $590,000

Hence remaining financing needed with long liabilities= $679,610 - $590,000= $89,610